OpenStudy (anonymous):

The area of a triangle is 16 square units. The base of the triangle is x+4. The height of the triangle is x. Find x.

OpenStudy (zzr0ck3r):

area of a triangle is \(A=\frac{1}{2}bh \) you have \(A=16, b=(x+4), h = x\).

OpenStudy (zzr0ck3r):

plugging in we get \(16=\frac{1}{2}(x+4)x=\frac{1}{2}x^2+2x\) can you solve \(\frac{1}{3}x^2+2x=16\)?

OpenStudy (anonymous):

is 1/2 of x^2 just x?